# Probing displaced (dark)photons from low reheating freeze-in at the LHC

### v1

@@ 1 Introduction
The early history of the Universe and the nature of dark matter (DM) remain deeply mysterious. Observationally, Big Bang Nucleosynthesis (BBN) represents the earliest epoch that can be directly probed, thanks to precise measurements of the primordial abundances. These observations offer a robust window into the thermal history of the Universe at temperatures of a few MeV Olive:1989xf ; Copi:1994ev . However, earlier stages remain largely unconstrained, allowing for the possibility of non-standard cosmological scenarios Hannestad:2004px ; Allahverdi:2020bys . As a result, many ideas have been proposed that modify the standard picture between the end of inflation and BBN. In particular, DM production in such non-standard cosmologies can exhibit features that differ significantly from traditional expectations. Examples include an early matter-dominated era, and scenarios with a low reheating temperature $T_{\rm RH}$ — defined as the temperature at which the post-inflationary Universe transitions into the standard hot Big Bang phase — among many other possibilities Giudice:2000ex ; Fornengo:2002db ; Pallis:2004yy ; Gelmini:2006pw ; Yaguna:2011ei ; Drees:2017iod ; Bernal:2018kcw ; Arias:2019uol ; Silva-Malpartida:2023yks ; Silva-Malpartida:2024emu .
@@
On another front, dark matter (DM) models in which the dark sector is secluded from the Standard Model (SM) and interacts only via mediator particles — through so-called portals — have attracted considerable attention. The freeze-in mechanism Hall:2009bx is a commonly invoked framework to generate such secluded dark sectors, characterized by the fact that DM never reaches thermal equilibrium with the visible sector. Traditionally, this non-thermal production has been studied under the assumption that the maximum temperature involved, $T_{\rm RH}$, lies well above all relevant mass scales in both the SM and the dark sector. As a result, extremely small couplings between the two sectors are typically required, making their experimental detection highly challenging.
@@
However, in recent years, the freeze-in production of DM at low $T_{\rm RH}$ has gained significant attention. This interest is driven by the realization that such cosmological scenarios are not only viable but also open new opportunities for laboratory tests. If $T_{\rm RH}$ lies below the masses of some dark-sector states, the corresponding coupling constants must increase to compensate for the Boltzmann suppression in the production rate. As a result, these models are testable not only in colliders but also in direct detection experiments Roszkowski:2014lga ; Cosme:2023xpa ; Koivunen:2024vhr ; Arcadi:2024wwg ; Boddy:2024vgt ; Arcadi:2024obp ; Lebedev:2024mbj ; Bernal:2024ndy ; Lee:2024wes ; Belanger:2024yoj ; Arias:2025nub ; Borah:2025ema ; Mondal:2025awq ; Khan:2025keb ; Bernal:2025qkj . Remarkably, reheating temperatures as low as $\mathcal{O}(4~{}\mathrm{MeV})$ remain consistent with BBN Giudice:2000ex ; Hannestad:2004px ; Allahverdi:2020bys , making them cosmologically well motivated Lebedev:2022cic .
@@
On the other hand, experimental collaborations at colliders have made remarkable efforts to improve their sensitivities to long-lived particles (LLPs), that is, particles with a lifetime that is long enough for them to travel a macroscopic distances before decaying Curtin:2018mvb ; Lee:2018pag ; Alimena:2019zri . This can lead to a series of non-standard signatures, such as displaced vertices, disappearing tracks, emerging jets and displaced photons. In particular, it is possible for the Higgs boson itself to decay into such states (see e.g. Strassler:2006im ; Strassler:2006ri ; Curtin:2013fra ; Cui:2014twa ; Craig:2015pha ; Gago:2015vma ; Caputo:2017pit ; Jones-Perez:2019plk ).
@@
Among the scenarios mentioned above, we highlight a case where the LLP can produce displaced visible photons in association with invisible particles, resulting in a non-pointing photon plus missing transverse energy signal at the LHC. This was first studied in the context of gauge mediated supersymmetry breaking models, with the lightest neutralino decaying into a photon and a gravitino ATLAS:2013etx ; ATLAS:2014kbb . Then, it was found it could probe the decays of heavy neutral leptons and DM with dipole interactions Primulando:2015lfa ; Duarte:2016caz ; Delgado:2022fea . However, the sensitivity reach of the search for non-pointing photons coming from the decays of pair-produced LLPs from Higgs bosons, with a dedicated lepton trigger ATLAS:2022vhr , opened the possibility to constrain this kind of exotic Higgs decays, and place stringent bounds on the Higgs interaction with the LLPs Duarte:2023tdw . These studies have been complemented by several proposals to search for photons from LLP decays in future facilities Barducci:2022gdv ; Barducci:2024nvd ; Bertuzzo:2024eds , or with new proposed search strategies Beltran:2024twr ; Kim:2025tuz .
@@
Interestingly, it has been shown that models featuring LLPs can also provide DM candidates, produced via freeze-in with low $T_{\rm RH}$, where the LLP acts as a mediator Co:2015pka ; Belanger:2018sti ; Calibbi:2021fld ; Bernal:2025qkj . Specifically, it was shown that searches for LLPs could place significant constraints on such models. Nevertheless, the possibility of having the mediator coupling to both photons and DM was not considered in detail in any of these previous studies. Thus, in this work we intend to bridge the gap, presenting a DM model with the relevant couplings, capable of generating the observed relic abundance, and demonstrating the capacity of displaced photons searches for applying bounds on the corresponding parameter space.
@@ 2 Model
In this work, we consider a dark sector consisting of a new real pseudo-scalar field $\phi$ and a dark vector boson $A^{\prime}_{\mu}$, the latter arising from a local $U^{\prime}(1)$ symmetry. Both fields are assumed to be massive, with the mass of $A^{\prime}_{\mu}$ generated either via a Stueckelberg mechanism or through a dark Higgs mechanism. [^1] For an explicit realization of a dark Higgs mechanism generating a mass for a $\mathbb{Z}_{2}$-odd $A^{\prime}_{\mu}$, see e.g. Farzan:2012hh . We impose a $\mathbb{Z}_{2}$ symmetry under which the dark fields $A^{\prime}_{\mu}$ and $\phi$ are odd, while all SM fields are even. This symmetry plays a crucial role in shaping the structure of the dark sector: it constrains the allowed couplings between the dark fields and prevents some direct interactions with the SM fields. In particular, it forbids kinetic mixing between the new $U^{\prime}(1)$ gauge boson and the SM hypercharge gauge boson. As a result, the renormalizable part of the Lagrangian is extended by:
@@
where $g_{D}$ is a dimensionful parameter, related to a high-energy scale, $g_{D}=1/\Lambda$, and $\tilde{B}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\lambda}B_{\rho\lambda}$ is the dual field strength of $U(1)_{Y}$. Considering that the model has four free parameters $(m_{\gamma^{\prime}},\,m_{\phi},\,\lambda_{HS},\,g_{D})$, in the following we identify the relevant parameter space based on the viability of having $\phi$ as a potential LLP.
@@ 2.1 Target parameter space
As mentioned in the introduction, we are interested in exploring the parameter space region of the model that can be accessed in collider experiments: more specifically, we want to reinterpret the results from displaced photon searches already carried out at the LHC  CMS:2012bbi ; ATLAS:2013etx ; ATLAS:2014kbb ; CMS:2019zxa ; ATLAS:2022vhr in terms of the relevant parameters of our model.
Indeed, the most recent search in ATLAS ATLAS:2022vhr looks for neutral long-lived particles produced from Higgs boson exotic decays, which eventually disintegrate into photons and missing energy. In our context, if $m_{\phi}<m_{h}/2$, the Higgs can decay into $\phi$ pairs via the Higgs portal, that is, the last term in eq. (1) with $\lambda_{HS}$ coupling. Furthermore, if $m_{\gamma^{\prime}}<m_{\phi}$, the pseudo-scalar becomes unstable, and decays into a photon and a dark photon via the dark photon portal, eq. (3), with $g_{D}$ coupling. This gives us the collider signal $pp\to h\to\phi\phi$, with each pseudo-scalar decaying as $\phi\to\gamma~{}\gamma^{\prime}$, which is exactly the signature probed by ATLAS:2022vhr . Therefore, in this work we will assume the hierarchy $m_{\gamma^{\prime}}<m_{\phi}<m_{h}/2$, with the dark photon $\gamma^{\prime}$ playing the role of stable DM. [^2] Notice that the inverse hierarchy $m_{\phi}<m_{\gamma^{\prime}}$ is a completely viable dark matter scenario (e.g. see Arias:2025nub ), with the $\phi$ being the DM candidate and $\gamma^{\prime}$ being the unstable dark state.
For displaced photon searches to be sensitive to our model, we need the scalar to travel enough before decaying: 1 cm $\lesssim c\,\tau_{\phi}\lesssim 100$ m. In addition, the displaced photon in the search must satisfy a cut on its transverse momentum, in fact, the search shows its maximum sensitivity when the particles producing the missing energy — in our case, the dark photons — are light enough to have the hardest possible visible photons. Based on previous work recasting this search Duarte:2023tdw , we focus on relatively light DM, with $m_{\gamma^{\prime}}\lesssim 5$ GeV.
@@
[…]
@@ 3 Dark matter production and cosmology
For our purposes, we assume the maximum temperature at which the dark sector particles are produced is the so-called reheating temperature $T_{\rm RH}$. [^3] Reheating is not expected to be truly an instantaneous process, therefore, the maximum temperature can be higher than $T_{\rm RH}$, see e.g. Giudice:2000ex ; Cosme:2024ndc . Even when other production processes may be present before reheating, we assume that before this period there is no population of $\gamma^{\prime}$ nor $\phi$. We also assume an instantaneous reheating epoch, as the SM sector is expected to thermalize rapidly due to its internal gauge couplings, after which the SM scattering and decays produce the dark sector from $T_{\rm RH}$ to lower temperatures.
Furthermore, we focus on scenarios with low reheating temperatures such that $T_{\rm RH}$ is comparable to or smaller than the new mass scales $m_{\gamma^{\prime}}$ and $m_{\phi}$. In this regime, Boltzmann suppression efficiently reduces the production rates, enabling the correct relic abundance to be achieved via freeze-in. This consideration proves of high importance since, as shown in previous works, the UV-freeze-in mediated by $g_{D}$, with high reheating temperatures in the parameter space identified at the end of section 2.1, typically leads to DM overproduction (see e.g. figure 3 of Arias:2025nub ). An alternative possibility is a fully thermal scenario DiazSaez:2024dzx ; however, it has been shown that this also leads to a significant overabundance in our target parameter space. In light of this, we explore the freeze-in production of dark states in the regime $m_{\gamma^{\prime}}\lesssim T_{\rm RH}\ll m_{\phi}$. Specifically, as shown later, to fulfill the correct relic abundance in the parameter space of our interest, we must have $T_{\rm RH}\gtrsim\mathcal{O}(1\,\rm{GeV})$.
@@
mediated by $\lambda_{HS}$ and $g_{D}$, respectively. Once dark states acquire non-negligible abundances, processes such as $\phi\leftrightarrow\gamma^{\prime}\gamma$ may impact the yield of the population already produced of $\gamma^{\prime}$ (see figure 2). [^4] We do not consider processes such as $t\bar{t}\rightarrow\gamma\textquoteright\phi,\,h\rightarrow\phi\phi,\,W^{+}W^{-}\rightarrow\gamma\textquoteright\phi,\,\phi h\rightarrow\gamma\textquoteright\gamma,\,f\phi\rightarrow f\gamma\textquoteright$ ’, and $\gamma\gamma\rightarrow\gamma\textquoteright\gamma\textquoteright\,(\phi\phi)$, as their amplitudes are subleading in our target parameter space. The first four processes are strongly Boltzmann suppressed, and using micrOMEGAs, we have verified that dark Primakov processes are also subdominant. The last process scales as $g_{D}^{4}$. In addition, because the Higgs portal $\lambda_{HS}$ can be relatively large, there are scenarios where $\phi$ may thermalize, thereby introducing new features to the yield evolution.

### v2

@@ 1 Introduction
The early history of the Universe and the nature of dark matter (DM) remain deeply mysterious. Observationally, Big Bang Nucleosynthesis (BBN) represents the earliest epoch that can be directly probed, thanks to precise measurements of the primordial abundances. These observations offer a robust window into the thermal history of the Universe at temperatures of a few MeV Olive et al. (1990); Copi et al. (1995). However, earlier stages remain largely unconstrained, allowing for the possibility of non-standard cosmological scenarios Kawasaki et al. (2000); Hannestad (2004); Hasegawa et al. (2019); Allahverdi and others (2020). As a result, many ideas have been proposed that modify the standard picture between the end of inflation and BBN. In particular, DM production in such non-standard cosmologies can exhibit features that differ significantly from traditional expectations. Examples include an early matter-dominated era, and scenarios with a low reheating temperature $T_{\rm RH}$ — defined as the temperature at which the post-inflationary Universe transitions into the standard hot Big Bang phase — among many other possibilities Giudice et al. (2001); Fornengo et al. (2003); Pallis (2004); Gelmini and Gondolo (2006); Yaguna (2012); Drees and Hajkarim (2018); Bernal et al. (2019); Arias et al. (2019); Silva-Malpartida et al. (2023, 2025).
@@
On another front, dark matter (DM) models in which the dark sector is secluded from the Standard Model (SM) and interacts only via mediator particles — through so-called portals — have attracted considerable attention. The freeze-in mechanism Hall et al. (2010) is a commonly invoked framework to generate such secluded dark sectors, characterized by the fact that DM never reaches thermal equilibrium with the visible sector. Traditionally, this non-thermal production has been studied under the assumption that the maximum temperature involved, $T_{\rm RH}$, lies well above all relevant mass scales in both the SM and the dark sector. As a result, extremely small couplings between the two sectors are typically required, making their experimental detection highly challenging.
@@
However, in recent years, the freeze-in production of DM at low $T_{\rm RH}$ has gained significant attention. This interest is driven by the realization that such cosmological scenarios are not only viable but also open new opportunities for laboratory tests. If $T_{\rm RH}$ lies below the masses of some dark-sector states, the corresponding coupling constants must increase to compensate for the Boltzmann suppression in the production rate. As a result, these models are testable not only in colliders but also in direct detection experiments Roszkowski et al. (2014); Bhattiprolu et al. (2023); Cosme et al. (2024a); Koivunen et al. (2024); Arcadi et al. (2024); Boddy et al. (2025); Arcadi et al. (2025); Barman et al. (2024, 2025); Lebedev et al. (2025); Bernal et al. (2025b); Lee et al. (2025); Bélanger et al. (2025); Arias et al. (2025); Borah et al. (2025); Mondal et al. (2025); Khan et al. (2025); Bernal et al. (2025a). Remarkably, reheating temperatures as low as $\mathcal{O}(4~\mathrm{MeV})$ remain consistent with BBN Kawasaki et al. (2000); Giudice et al. (2001); Hannestad (2004); Hasegawa et al. (2019); Allahverdi and others (2020), making them cosmologically well motivated Lebedev (2023).
@@
On the other hand, experimental collaborations at colliders have made remarkable efforts to improve their sensitivities to long-lived particles (LLPs), that is, particles with a lifetime that is long enough for them to travel a macroscopic distances before decaying Curtin and others (2019); Lee et al. (2019); Alimena and others (2020). This can lead to a series of non-standard signatures, such as displaced vertices, disappearing tracks, emerging jets and displaced photons. In particular, it is possible for the Higgs boson itself to decay into such states (see e.g. Strassler and Zurek (2007, 2008); Curtin and others (2014); Cui and Shuve (2015); Craig et al. (2015); Gago et al. (2015); Caputo et al. (2017); Jones-Pérez et al. (2020)).
@@
Among the scenarios mentioned above, we highlight a case where the LLP can produce displaced visible photons in association with invisible particles, resulting in a non-pointing photon plus missing transverse energy signal at the LHC. This was first studied in the context of gauge mediated supersymmetry breaking models, with the lightest neutralino decaying into a photon and a gravitino Aad and others (2013, 2014). Then, it was found it could probe the decays of heavy neutral leptons and DM with dipole interactions Primulando et al. (2015); Duarte et al. (2018); Delgado et al. (2022). However, the sensitivity reach of the search for non-pointing photons coming from the decays of pair-produced LLPs from Higgs bosons, with a dedicated lepton trigger Aad and others (2023b), opened the possibility to constrain this kind of exotic Higgs decays, and place stringent bounds on the Higgs interaction with the LLPs Duarte et al. (2024). These studies have been complemented by several proposals to search for photons from LLP decays in future facilities Barducci et al. (2023, 2024); Bertuzzo and Frigerio (2025), or with new proposed search strategies Beltrán et al. (2024); Kim et al. (2025).
@@
Interestingly, it has been shown that models featuring LLPs can also provide DM candidates, produced via freeze-in with low $T_{\rm RH}$, where the LLP acts as a mediator Co et al. (2015); Bélanger and others (2019); Calibbi et al. (2021); Bernal et al. (2025a). Specifically, it was shown that LHC searches for LLPs could place significant constraints on such models. Nevertheless, the possibility of having an LLP mediator coupling to both photons and DM without reaching thermalization was not considered in detail in any of these previous studies. Thus, in this work we intend to bridge the gap, presenting a DM model with the relevant couplings, capable of generating the observed relic abundance, and demonstrating the capacity of displaced photons searches for applying bounds on the corresponding parameter space.
@@ 2 Model
In this work, we consider a dark sector consisting of a new real pseudo-scalar field $\phi$ and a dark vector boson $A^{\prime}_{\mu}$, the latter arising from a local $U^{\prime}(1)$ symmetry. Both fields are assumed to be massive, with the mass of $A^{\prime}_{\mu}$ generated either via a Stueckelberg mechanism or through a dark Higgs mechanism. [^1] For an explicit realization of a dark Higgs mechanism generating a mass for a $\mathbb{Z}_{2}$-odd $A^{\prime}_{\mu}$, see e.g. Farzan and Akbarieh (2012). We impose a $\mathbb{Z}_{2}$ symmetry under which the dark fields $A^{\prime}_{\mu}$ and $\phi$ are odd, while all SM fields are even. This symmetry plays a crucial role in shaping the structure of the dark sector: it constrains the allowed couplings between the dark fields and prevents some direct interactions with the SM fields. In particular, it forbids kinetic mixing between the new $U^{\prime}(1)$ gauge boson and the SM hypercharge gauge boson. As a result, the renormalizable part of the Lagrangian is extended by:
@@
where $g_{D}$ is a dimensionful parameter, related to a high-energy scale, $g_{D}=1/\Lambda$, and $\tilde{B}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\lambda}B_{\rho\lambda}$ is the dual field strength of $U(1)_{Y}$. Considering that the model has four free parameters [^2] From the collider perspective, $\lambda_{\phi}$ does not play a relevant role, so we do not define any specific target range. However, it may have an important impact on the cosmological evolution of the particle content; we therefore comment on it in the next section. $(m_{\gamma^{\prime}},\,m_{\phi},\,\lambda_{HS},\,g_{D})$, in the following we identify the relevant parameter space based on the viability of having $\phi$ as a potential LLP.
@@ 2.1 Target parameter space
As mentioned in the introduction, we are interested in exploring the parameter space region of the model that can be accessed in collider experiments: more specifically, we want to reinterpret the results from displaced photon searches already carried out at the LHC  Chatrchyan and others (2013); Aad and others (2013, 2014); Sirunyan and others (2019b); Aad and others (2023b) in terms of the relevant parameters of our model.
Indeed, the most recent search in ATLAS Aad and others (2023b) looks for neutral long-lived particles produced from Higgs boson exotic decays, which eventually disintegrate into photons and missing energy. In our context, if $m_{\phi}<m_{h}/2$, the Higgs can decay into $\phi$ pairs via the Higgs portal, that is, the last term in eq. (1) with $\lambda_{HS}$ coupling. Furthermore, if $m_{\gamma^{\prime}}<m_{\phi}$, the pseudo-scalar becomes unstable, and decays into a photon and a dark photon via the dark photon portal, eq. (3), with $g_{D}$ coupling. This gives us the collider signal $pp\to h\to\phi\phi$, with each pseudo-scalar decaying as $\phi\to\gamma~\gamma^{\prime}$, which is exactly the signature probed by Aad and others (2023b). Therefore, in this work we will assume the hierarchy $m_{\gamma^{\prime}}<m_{\phi}<m_{h}/2$, with the dark photon $\gamma^{\prime}$ playing the role of stable DM. [^3] Notice that the inverse hierarchy $m_{\phi}<m_{\gamma^{\prime}}$ is a completely viable dark matter scenario (e.g. see Arias et al. (2025)), with the $\phi$ being the DM candidate and $\gamma^{\prime}$ being the unstable dark state.
For displaced photon searches to be sensitive to our model, we need the scalar to travel enough before decaying: 1 cm $\lesssim c\,\tau_{\phi}\lesssim 100$ m. In addition, the displaced photon in the search must satisfy a cut on its transverse momentum, in fact, the search shows its maximum sensitivity when the particles producing the missing energy — in our case, the dark photons — are light enough to have the hardest possible visible photons. Based on previous work recasting this search Duarte et al. (2024), we focus on relatively light DM, with $m_{\gamma^{\prime}}\lesssim 5$ GeV.
@@
where $\Delta m^{2}_{\phi\gamma^{\prime}}=m^{2}_{\phi}-m^{2}_{\gamma^{\prime}}$ is compared to a benchmark value for $m_{\gamma^{\prime}}=10$ GeV and $m_{\phi}=60$ GeV. In order to illustrate this result, we show in figure 1 values of $c\,\tau_{\phi}$ in the $g_{D}-m_{\phi}$ plane. The curves correspond to $c\,\tau_{\phi}=0.01,\,1,\,100$  m, with vanishing values of $m_{\gamma^{\prime}}$ (solid lines), or equal to $m_{\phi}-15$ GeV (dashed lines). As discussed in section 4, the displaced photon search is most sensitive to LLP displacements around $c\,\tau_{\phi}\approx 1$ m. Therefore, the figure suggests that the relevant parameter space lies within $g_{D}\sim 10^{-10}$ –- $10^{-9}$ GeV-1. Notice that, due to the $Z\phi\gamma^{\prime}$ interaction induced by the dark photon portal in eq. (3), the $g_{D}$ coupling is constrained by LEP measurements of the $Z$ boson invisible width and $Z\to\gamma+{\rm inv.}$, depending on whether the pseudo-scalar decays outside the detector or not. In Jodłowski (2024) the bound from LEP (in the $m_{\gamma^{\prime}}\ll m_{\phi}\lesssim 60$ GeV limit) is found to require $g_{D}\lesssim 10^{-4}$ GeV-1, far above our target scenario.
@@
Regarding the values of $\lambda_{HS}$, some regions are subject to existing constraints from Higgs data, which we will discuss in section 4. Nonetheless, we include this broader range in section 3, for illustrative purposes. Concerning the coupling $g_{D}$ our target parameter space remains unexplored in the mass range we propose to scan. On the other hand, stronger constraints apply in the sub-MeV range (e.g. Arias et al. (2021); Hook et al. (2024)).
@@ 3 Dark matter production and cosmology
For our purposes, we assume the maximum temperature at which the dark sector particles are produced is the so-called reheating temperature $T_{\rm RH}$. [^4] Reheating is not expected to be truly an instantaneous process, therefore, the maximum temperature can be higher than $T_{\rm RH}$, see e.g. Giudice et al. (2001); Cosme et al. (2024b). Even when other production processes may be present before reheating, we assume that before this period there is no population of $\gamma^{\prime}$ nor $\phi$. We also assume an instantaneous reheating epoch, as the SM sector is expected to thermalize rapidly due to its internal gauge couplings, after which the SM scattering and decays produce the dark sector from $T_{\rm RH}$ to lower temperatures.
Furthermore, we focus on scenarios with low reheating temperatures such that $T_{\rm RH}$ is comparable to or smaller than the new mass scales $m_{\gamma^{\prime}}$ and $m_{\phi}$. In this regime, Boltzmann suppression efficiently reduces the production rates, enabling the correct relic abundance to be achieved via freeze-in. This consideration proves of high importance since, as shown in previous works, the UV-freeze-in mediated by $g_{D}$, with high reheating temperatures in the parameter space identified at the end of section 2.1, typically leads to DM overproduction (see e.g. figure 3 of Arias et al. (2025)). An alternative possibility is a fully thermal scenario Díaz Sáez (2025); however, it has been shown that this also leads to a significant overabundance in our target parameter space. In light of this, we explore the freeze-in production of dark states in the regime $m_{\gamma^{\prime}}\lesssim T_{\rm RH}\ll m_{\phi}$. Specifically, as shown later, to fulfill the correct relic abundance in the parameter space of our interest, we must have $T_{\rm RH}\gtrsim\mathcal{O}(1\,\rm{GeV})$.
@@
mediated by $\lambda_{HS}$ and $g_{D}$, respectively. Once dark states acquire non-negligible abundances, processes such as $\phi\leftrightarrow\gamma^{\prime}\gamma$ may impact the yield of the population already produced of $\gamma^{\prime}$ (see figure 2). [^5] We do not consider the processes $t\bar{t}\rightarrow\gamma^{\prime}\phi,\,h\rightarrow\phi\phi,\,hh\rightarrow\phi\phi,\,W^{+}W^{-}\rightarrow\gamma^{\prime}\phi,\,\phi h\rightarrow\gamma^{\prime}\gamma,\,f\phi\rightarrow f\gamma^{\prime}$, and $\gamma\gamma\rightarrow\gamma^{\prime}\gamma^{\prime}\,(\phi\phi)$, as their amplitudes are subleading in our target parameter space. The first five processes are strongly Boltzmann suppressed, and using micrOMEGAs, we have verified that dark Primakov processes are also subdominant. The last process scales as $g_{D}^{4}$. In addition, because the Higgs portal $\lambda_{HS}$ can be relatively large, there are scenarios where $\phi$ may thermalize, thereby introducing new features to the yield evolution. To avoid self-production and cannibalization processes, in the following analysis, we assume that $\lambda_{\phi}$ is small enough not to affect the relic abundance through dark freeze-out (see e.g., Bernal and Chu (2016); Heikinheimo et al. (2017); Arcadi et al. (2019); Bernal (2020) for related discussions).